I am supposed to solve systems using elimination but i do not understand especially when it doesnt have a nber for example -4x-15y=-17 -x+5y=-13 what do i do ?
Answer by Alan3354(69443)

-4x-15y=-17
-x + 5y=-13
=============
Doesn't have a number?
Multiply the 2nd eqn by 3
-4x-15y=-17
-3x+15y=-39
------------ Add
-7x = -56
x = 8
-------
-x + 5y=-13
-8 + 5y = -13
y = -1